OpenStudy (anonymous):

-2x - 18 = 6(1 - 2x) + 4x x =

OpenStudy (andriod09):

Your trying to find "x", I presume, so this is the equation, correct? \[-2x-18=6(1-2x)+4x\]

OpenStudy (jack1):

2...?

OpenStudy (nurali):

-2x - 18 = 6(1 - 2x) + 4x -2x-18=6-12x+4x -2x-18=6-8x -2x+8x=6+18 6x=24 x=24/6 x=4